Should print ONTWERP only at the bottom each page.

\usepackage[none,bottom]{draftcopy}

Test5: This is a test page 1

This is a testpage 1 Test5: This is a test page 2 This is a testpage 2 Test5: This is a test page 3 This is a testpage 3 Test5: This is a test page 4 This is a testpage 4